# First moments of ${\rm GL}(3)\times{\rm GL}(2)$ and ${\rm GL}(2)$ $L$-functions and their applications to subconvexity

### v1

@@ Abstract.
Let $F$ be a self-dual Hecke-Maaß form for ${\rm GL}(3)$ underlying the symmetric square lift of a ${\rm GL}(2)$-newform of square-free level and trivial nebentypus. In this paper, we are interested in the first moments of the central values of $\rm GL(3)\times GL(2)$ $L$-functions and $\rm GL(2)$ $L$-functions. As a result, we obtain an estimate for the first moment for $L(1/2,F\otimes f)$ over a family, where $F$ is of the level $q^{2}$, and $f\in\mathcal{B}^{\ast}_{k}(M)$ for any primes $q,M\geq 2$ such that $(q,M)=1$. We prove the subconvex bound for $L(1/2,F\otimes f)$ involving the levels aspects simultaneously in the range $M^{5/32+\varepsilon}\leq q\leq M^{7/32-\varepsilon}$ and $M>q^{\delta}$ for any $\varepsilon,\delta>0$ for the first time. Moreover, we present a first moment estimate for $L(1/2+it,F\otimes f)$ over $K\leq k\leq 2K$, with $K$ being a large number. As an application, whenever $k^{10/9+\varepsilon}\leq|t|\leq k^{2-\varepsilon}$, we obtain a non-trivial estimate for the first moment of $L(1/2,f)L(1/2+it,F\otimes f)$ in the weight $k$ and $t$ aspects.
@@ 1.1. Overview
Let $N_{1},N_{2}\geq 1$ be two integers. Let $k$ be an even integer and $t\geq 1$. For any normalized ${{\rm GL}}(3)$-Maaß form $F$ of level $N_{1}$ and ${{\rm GL}}(2)$-Maaß form $f$ of level $N_{2}$ and weight $k$, there holds that $L(1/2+it,F\otimes f)\ll_{\varepsilon}(N_{1}N^{3/2}_{2}k^{3}t^{3})^{1/2-\varepsilon}$ for some absolute constant $\varepsilon>0$.
@@
    when $N_{1},N_{2}$ are both co-prime primes. Their bound is subconvex in the range $N^{1/2+\varepsilon}_{2}<N_{1}<N^{3/2-\varepsilon}_{2}$.
@@ 1.2. Main results
In this paper, we are more concerned about the self-dual ${\rm GL}(3)\times{\rm GL}(2)$ $L$-functions, which have important implications to equidistribution problems of $L^{q}$-mass of cusp forms (as $q\rightarrow\infty$) and Heegner points on Shimura curves as well as the Quantum Unique Ergodicity (QUE) conjecture (see, for example, [4, 14, 3, 35, 34, 13, 6, 42]). For instance, the well-known Watson’s formula implies
@@
for any ${\rm GL}(2)$-newform $g\in\mathcal{B}^{\ast}_{k}(p)$ of large fixed weight $k$ and prime level $p$. Notice that the level aspect subconvexity for ${\rm GL}(3)$ forms is still wide-open for specialists, albeit the notable successes of subconvexity for ${\rm GL}(3)$ $L$-functions in various aspects over the past decade. One of main objectives of the present paper is to address the subconvexity in the levels aspects for ${\rm GL}(3)$ and ${\rm GL}(2)$ forms simultaneously [^2] One might wander whether the method of Kumar, Munshi and Singh can be adpated to obtain subconvexity for $L(1/2,\text{sym}^{2}g\otimes f)$ with $F=\text{sym}^{2}g$ being a symmetric square lift of a ${\rm GL}(2)$-newform $g$, say. Unfortunately, this seems impracticable. The underlying reason is that the modulus $\mathcal{Q}$ in equipping with $\delta$-symbol method is usually large relative to the level of $F$; see [29, Section 2-2B]. This gives rise to a deadlock since of an inevitable gap for ${\rm GL}(3)$-Voronoĭ formulae which we will explain after a while.. To be specific, we prove the following results:
@@
Let $q,M\geq 2$ be two primes satisfying that $(q,M)=1$ and $q<M^{3/8}$. Let $F$ be a Hecke-Maaß form for the congruence subgroup $\Gamma_{0}(q^{2})$ obtained as the symmetric square lift of a ${\rm GL}(2)$-newform of level $q$ with trivial nebentypus. Define the set
$\mathscr{P}=\{p,p^{2}:p\text{ is a prime },(p,qM)=1,L\leq p\leq 2L\}$
@@
$\frac{q^{2+\varepsilon}}{{\sqrt{M}}}+\frac{\sqrt{M}}{q^{2-\varepsilon}}\leq L\leq\min{\left\{q^{2-\varepsilon}\sqrt{M},\frac{M^{1-\varepsilon}}{q^{2}}\right\}}.$
Then, for any $\ell\in\mathscr{P}$, we have
@@
Observe that under the self-duality for $\rm GL(3)$ forms, the non-negativity for the central values of the ${\rm GL}(3)\times{\rm GL}(2)$ $L$-functions is guaranteed by Lapid’s method [30]; see also [32, Remark 3]. As an application of Theorem 1.2, we are thus able to obtain the following subconvex bound:
@@
$\frac{q^{2+\varepsilon}}{{\sqrt{M}}}+\frac{\sqrt{M}}{q^{2-\varepsilon}}+\frac{M^{1/4+\varepsilon}}{q}\leq L\leq\min{\left\{M^{3/16-\varepsilon},\frac{M^{3/8-\varepsilon}}{q}\right\}},$ (1.3)
@@
$L(1/2,F\otimes f)\ll\mathcal{X}^{1/2+\varepsilon}\left(\frac{M^{1/4}}{qL}+\frac{\sqrt{L}}{M^{1/4}}+\frac{Lq}{M^{3/8}}+\frac{L^{2}}{M^{3/8}}+\frac{L^{7/4}q^{5/2}}{M^{7/8}}+\frac{1}{\sqrt{M}}\right).$ (1.4)
This establishes the subconvexity for any $q,M$ with $M^{5/32+\varepsilon}\leq q\leq M^{7/32-\varepsilon}$ and $M>q^{\delta}$ for any $\varepsilon,\delta>0$, and should be the first instance for subconvexity in the levels aspects for ${\rm GL}(3)\times{\rm GL}(2)$. Taking a little effort, it is not hard to attain the subconvexity bounds for any $q,M$ varying in this range $M^{5/32+\varepsilon}\leq q\leq M^{7/32-\varepsilon}$. For instance, when $q\asymp M^{3/16}$, we have
$L(1/2,F\otimes f)\ll\mathcal{X}^{1/2-1/30+\varepsilon}.$
In this paper, another motivation is to investigate the first moments of $L(1/2+it,F\otimes f)$ and $L(1/2,f)L(1/2+it,F\otimes f)$ in the weight $k$ and $t$ aspects over families. We continue to explore the first moment method by adding a ‘harmonic’ average of the weight $k$, with $K<k\leq 2K$. As a result, we are able to establish
@@
Let $F(z)$ be a normalized Hecke-Maaß form on ${{\rm GL}}(3)$ (not necessarily self-dual). Let $k$ be an even integer. Let $K$ be a sufficiently large parameter and $K^{5/9+\varepsilon}<|t|\ll K^{2-\varepsilon}$. Let $W$ be a smooth function compactly supported on $[1/2,5/2]$ with bounded derivatives. Then, for any integer $\ell\geq 1$ satisfying that $K^{1+\varepsilon}/|t|^{3}\leq\ell\leq|t|^{3-\varepsilon}/K^{7/3}$, and any ${{\rm GL}}(2)$-newform $f$ of weight $k$ and full level, we have
$\sum_{k\equiv 0\bmod 2}W{\left(\frac{k-1}{K}\right)}\sum_{f\in\mathcal{B}^{\ast}_{k}}\omega^{-1}_{f}\lambda_{f}(\ell)L(1/2+it,F\otimes f)$
$\hskip 122.34692pt\ll(K|t|)^{\varepsilon}\left(\ell^{7/16}K^{27/16}|t|^{27/16}+\frac{\ell^{5/8}|t|^{23/8}}{K^{1/8}}\right).$ (1.5)
Consequently, as a by-product we can obtain a cute result as follows
@@
Let the forms $F,f$, the parameters $k,K$ and the function $W$ be as in Theorem 1.4. Then, for any ${{\rm GL}}(2)$-newform $f$ of weight $k$ and full level, and $K^{10/9+\varepsilon}\leq|t|\leq K^{2-\varepsilon}$, we have
$\sum_{k\equiv 0\bmod 2}W{\left(\frac{k-1}{K}\right)}\sum_{f\in\mathcal{B}^{\ast}_{k}}\omega^{-1}_{f}L(1/2,f)L(1/2+it,F\otimes f)\ll(K|t)^{\varepsilon}K^{21/8}|t|^{27/16}.$ (1.6)
This bound is already able to beat the convexity barrier in the weight $k$ and $t$ aspects, when they are varying in the range $k^{10/9+\varepsilon}\leq t\leq k^{2-\varepsilon}$ with $k$ being sufficiently large. By the convexity, the left-hand side of (1.6) is $\ll k^{3+\varepsilon}|t|^{3/2}$; while, the right-hand side shows a smaller magnitude by $\ll k^{21/8+\varepsilon}|t|^{27/16}$. Although one cannot expect the individual subconvexity bound for $L(1/2,f)L(1/2+it,F\otimes f)$, the estimate (1.4), however, reflects the effectiveness by further averaging over the parameter $k\sim K$, when it comes to investagting the subconvexity problem involving the weight aspect. We also refer the readers to the works of Luo [36] and Khan [25] for relevant heuristics. It is, on the other hand, believed that our method can be adapted to produce simultaneous non-vanishing of ${\rm GL}(3)\times{\rm GL}(2)$ and ${\rm GL}(2)$ $L$-functions at the central point, and even subconvexity for $L(1/2,F\otimes f)$ or $L(1/2,f)L(1/2,F\otimes f)$ in the weight and the level (of $F$ or $f$) aspects. We are now working on this direction, and expect to report the progress in the future article.
The initial manoeuvre of our proofs of main results follows the path pioneered by Xiaoqing Li and Blomer, who looked into the average of first moment of $\rm GL(3)\times GL(2)$ $L$-function over a family. The strategies we pursued in the paper, however, can be understood as the “Kuznetsov-Voronoĭ-Reciprocity-Voronoĭ-Bilinear form” in proving Theorem 1.2 and “Kuznetsov-Voronoĭ-Exponential sums” in proving Theorem 1.4, respectively. In particular, for the former, this is reminiscent of an important work of Blomer and Khan, [5], in which they investigated the twisted moments of $\text{GL}(3)\times\text{GL}(2)$ and $\text{GL}(2)$ automorphic $L$-functions, and established the following reciprocity formula
@@
by the “Kuznetsov-Voronoĭ-Reciprocity-Voronoĭ-Kuznetsov” law. Among other things, one salient point is the application of the reciprocity law, which, as the Kloosterman sum of dimension one, keeps us going back to somewhere. However, when applying the Voronoĭ to the Kloosterman sum $S(n\ell,1;c)$ of dimension two, the inverse of the variable $\ell$, that is, the exponential factor like $S(n\overline{\ell},\sigma;c)$ for some less significant parameter $\sigma$, will be occurred; this will increase the dimensions of Kloosterman sums afterwards, if one were to resort to the Voronoĭ again. In this sense, we can barely expect the reciprocity formula for Hecke-Maaß forms of higher degrees $d\geq 4$ in the levels aspects. It seems that the possible symmetric relation like (1.7) for $\text{GL}(d)\times\text{GL}(2)$ involving the levels aspects is only for the $\text{GL}(3)\times\text{GL}(2)$ case (we refer the interested readers to Jana and Nunes’s recent work [23] for the genernlization to $\rm{PGL}_{d+1}\times\rm{PGL}_{d-1}$). In contrast to the $\text{GL}(3)\times\text{GL}(2)$ case, we notice that, for the $\text{GL}(2)\times\text{GL}(1)$ case, Conrey [9] proved the conductors aspects reciprocity formula
@@
Let us now elucidate two of main obstacles in present paper. It is remarkable that, by the current development of the Voronoĭ formulae for ${\rm GL}(3)$ in the level aspect, there is still a gap for arbitrary non-zero integer $c\in\mathbb{Z}$ (see Lemma 2.1 below). If the additive twist colludes with the level $N$ of the form $F$, namely the parameters $c,N$ satisfy that $(N/(N,c),c)\neq 1$. For instance, one takes $N=q^{2}$ and $c=c^{\prime}q$ with $(c^{\prime},q)=1$ for any prime $q$. The ${{\rm GL}}(3)$-Voronoĭ formula in this scase looks rather tricky. In principle, Corbett’s formula [8, Theorem 1.1] can cover this, but it requires a non-trivial analysis of the $p$-adic Bessel transforms, which, however, becomes more involved; see [8] for relevant details. We also refer the readers to the work of Booker, Milinovich and Ng [7], who worked out the Voronoĭ formula in the ${{\rm GL}}(2)$-case (if the additive twist colludes with the level). As an innovation in this paper, we are able to overcome this deadlock, which makes the hybrid subconvexity for ${\rm GL}(3)$ and ${\rm GL}(2)$ in the levels aspects come into being. The key point, however, is to invoke the reciprocity law as mentioned above. Indeed, with a suitable constraint for the parameter $\ell$ we are in the unramified case, when applying the Voronoĭ for the first time; then, one would be transferred to the ramified case after an implementation of the reciprocity law. This makes it possible to equip with the Voronoĭ again.
The second obstacle is the root number. It is well-known that the explicit determination of root numbers play a focal rôle when doing averages of central values of $L$-functions over different families. In this paper, we will give a detailed description of the root number $\varepsilon(F\otimes f)$ for any ${\rm GL}(3)$ Hecke-Maaß form $F$ with arbitrary level $N$ in the framework of the adelic interpretations; see §2.3. Part of descriptions in the symmetric square lift case has already been given by Holowinsky, Munshi and Qi [16]. As it can be seen later, the root number $\varepsilon(F\otimes f)$ depends only on the parameter $k$ and the form $F$. One of its factor $i^{k}$ will have a direct influence on the ‘harmonic’ averages of the Bessel function $J_{k-1}$ in considering the moments of $L$-values in the weight $k$ aspect; see Lemmas 2.8 and 2.9.
@@
Throughout the paper, $\varepsilon$ (resp. $\delta$) always denotes an arbitrarily small positive constant which may not be the same at each occurrence. $A\ll_{\varepsilon}B$ means that $|A|\ll X^{\varepsilon}|B|$, and $n\sim N$ means that $N<n\leq 2N$ for every $n\in\mathbb{Z}$. As usual, the symbols $\mathbb{R}$, $\mathbb{C}$, $\mathbb{Z}$ and $\mathbb{N}^{+}$ are respectively denote the real number field, the complex number field, the ring of integers and the ring of positive integers.

### v4

# First moments of ${{\mathrm{GL}}}(3)\times{{\mathrm{GL}}}(2)$ and ${{\mathrm{GL}}}(2)$ $L$-functions and their applications

@@ Abstract.
Let $F$ be a self-dual Hecke-Maaß form for ${\mathrm{GL}}(3)$ underlying the symmetric square lift of a ${\mathrm{GL}}(2)$-newform of square-free level and trivial nebentypus. In this paper, we are interested in the first moments of the central values of $\rm GL(3)\times GL(2)$ $L$-functions and $\rm GL(2)$ $L$-functions. As a result, we obtain an estimate for the first moment for $L(1/2,F\otimes f)$ in a family, where $F$ is of the level $q^{2}$, and $f\in\mathcal{B}^{\ast}_{k}(M)$ for any primes $q,M\geq 2$ such that $(q,M)=1$. We prove the subconvex bound for $L(1/2,F\otimes f)$ involving the levels aspects simultaneously in the range $M^{13/64+\varepsilon}\leq q\leq M^{11/40-\varepsilon}$ and $M>q^{\delta}$ for any $\varepsilon,\delta>0$ for the first time. Moreover, we further investigate the first moments of these $L$-functions in the weight $k$ aspect over $K\leq k\leq 2K$, with $K$ being a large number. As the results, we obtain a Lindelöf average bound for the first moment of $L(1/2,f)L(1/2,F\otimes f)$ of degree 8 and an asymptotic formula for the first moment of $L(1/2,F\otimes f)$ with an error term of $O(K^{-1/4+\varepsilon})$, respectively.
@@ 1. Introduction
Let $N_{1},N_{2}\geq 1$ be two integers. Let $k$ be an even integer and $t\geq 1$. For any normalized ${{\mathrm{GL}}}(3)$-Maaß form $F$ of level $N_{1}$ and ${{\mathrm{GL}}}(2)$-Maaß form $f$ of level $N_{2}$ and weight $k$, there holds that $L(1/2+it,F\otimes f)\ll_{\varepsilon}(N_{1}N^{3/2}_{2}k^{3}t^{3})^{1/2-\varepsilon}$ for some absolute constant $\varepsilon>0$.
@@
    when $N_{1},N_{2}$ are both co-prime primes. This hybrid bound is subconvex in the range $N^{1/2+\varepsilon}_{2}<N_{1}<N^{3/2-\varepsilon}_{2}$.
@@ 1.1. Main results
In this paper, we are more concerned about the self-dual ${\mathrm{GL}}(3)\times{\mathrm{GL}}(2)$ $L$-functions, which have important implications to equidistribution problems of $L^{q}$-mass of cusp forms (as $q\rightarrow\infty$) and Heegner points on Shimura curves as well as the Quantum Unique Ergodicity (QUE) conjecture (see, for example, [5, 16, 3, 38, 37, 15, 7, 44]). For instance, the well-known Watson’s formula implies that
@@
for any ${\mathrm{GL}}(2)$-newform $g\in\mathcal{B}^{\ast}_{k}(p)$ of large fixed weight $k$ and prime level $p$. Notice that the level aspect subconvexity for self-dual ${\mathrm{GL}}(3)$ forms is still wide-open for specialists, albeit the notable successes of subconvexity for ${\mathrm{GL}}(3)$ $L$-functions in various aspects over the past decade. One of main objectives of the present paper is to address the subconvexity in the levels aspects for ${\mathrm{GL}}(3)$ and ${\mathrm{GL}}(2)$ forms simultaneously [^1] One might wander whether the method of Kumar, Munshi and Singh can be adpated to obtain subconvexity for $L(1/2,\text{sym}^{2}g\otimes f)$ with $F=\text{sym}^{2}g$ being a symmetric square lift of a ${\mathrm{GL}}(2)$-newform $g$, say. Unfortunately, this seems impracticable. The underlying reason is that the modulus $\mathcal{Q}$ in equipping with $\delta$-symbol method is usually large relative to the level of $F$; see [32, Section 2-2B]. This gives rise to a deadlock since of an inevitable gap for ${\mathrm{GL}}(3)$-Voronoĭ formulae which we will explain after a while.. To be specific, we prove the following results:
@@
Let $q,M\geq 2$ be two primes satisfying that $(q,M)=1$ and $M^{1/8}<q<M^{3/8}$. Let $F$ be a Hecke-Maaß form for the congruence subgroup $\Gamma_{0}(q^{2})$ obtained as the symmetric square lift of a ${\mathrm{GL}}(2)$-newform of level $q$ with trivial nebentypus. Define the set
$\mathscr{P}_{L}=\{p:p\text{ is a prime },(p,qM)=1,L\leq p\leq 2L\}$ (1.2)
@@
$\frac{q^{1+\varepsilon}}{{M^{1/4}}}+\frac{{M}^{1/4}}{{q^{1-\varepsilon}}}\leq L\leq M^{1/8-\varepsilon}.$
Then, for any ={1,$\ell=\{1,\ell_{1}\ell_{2},\ell^{2}_{1}\ell^{2}_{2}\}$ with $\ell_{1},\ell_{2}\in\mathscr{P}_{L}$, we have
@@
Observe that under the self-duality for $\rm GL(3)$ forms, the non-negativity for the central values of the ${\mathrm{GL}}(3)\times{\mathrm{GL}}(2)$ $L$-functions is guaranteed by Lapid’s method [33]; see also [35, Remark 3]. As an application of Theorem 1.2, we are thus able to obtain the following subconvex bound:
@@
$\frac{q^{1+\varepsilon}}{{M^{1/4}}}+\frac{M^{(1+3\delta)/4+\varepsilon}}{q^{1-\delta}}\leq L\leq\min{\left\{M^{3/64-\varepsilon},\frac{M^{3/32-\varepsilon}}{q^{1/4}},\,\frac{{M}^{1/8}}{q^{5/14+\varepsilon}},\,\frac{M^{7/40-\varepsilon}}{q^{3/10}}\right\}},$ (1.4)
@@
$L(1/2,F\otimes f)\ll\mathcal{X}^{1/2+\varepsilon+\delta}\left(\frac{M^{1/4}}{qL}+\frac{L^{2}}{M^{1/4}}+\frac{L^{4}q}{M^{3/8}}+\frac{L^{8}}{M^{3/8}}+\frac{L^{7}q^{5/2}}{M^{7/8}}+\frac{1}{\sqrt{M}}\right).$ (1.5)
This establishes the subconvexity for any $q,M$ with $M^{13/64+\varepsilon}\leq q\leq M^{11/40-\varepsilon}$ and $M>q^{\delta}$ for any $\varepsilon,\delta>0$, and should be the first instance for the levels aspects subconvexity of self-dual ${\mathrm{GL}}(3)\times{\mathrm{GL}}(2)$ $L$-functions. Taking a little effort, it is not hard to attain the subconvexity bounds for any $q,M$ varying in this range. For instance, when $q\asymp M^{1/4}$, we have
$L(1/2,F\otimes f)\ll\mathcal{X}^{1/2-1/80+\varepsilon}$
and when $q\asymp M^{11/40-\varepsilon}$, we obtain
$L(1/2,F\otimes f)\ll\mathcal{X}^{1/2-3/82+\varepsilon}.$
In this paper, another motivation is to investigate the first moments of $L(1/2,F\otimes f)$ and $L(1/2,f)L(1/2,F\otimes f)$ in the weight aspect in families. We continue to explore the first moment method by adding a ‘harmonic’ average of the weight $k$, with $K<k\leq 2K$. As a result, we are able to establish the following Lindelöf average bound:
@@
Let $F(z)$ be a normalized self-dual Hecke-Maaß form on ${{\mathrm{GL}}}(3)$. Let $K$ be a sufficiently large parameter. Let $W$ be a smooth function compactly supported on $[1/2,5/2]$ with bounded derivatives. Then, we have
$\sum_{k\equiv 0\bmod 2}W{\left(\frac{k-1}{K}\right)}\sum_{f\in\mathcal{B}^{\ast}_{k}}\omega^{-1}_{f}L(1/2,f)L(1/2,F\otimes f)\ll K^{1+\varepsilon},$ (1.6)
where the implied constant depends only on the constant $\varepsilon$, the form $F$ as well as the level of the form $f$.
In addition, we are able to obtain the following asymptotic formula for the ${\mathrm{GL}}(3)\times{\mathrm{GL}}(2)$ $L$-functions in the weight aspect with a power-saving error term:
Theorem 1.5.
Keep the notation as in Theorem 1.4. Then, we have
$\sum_{k\equiv 0\bmod 2}W{\left(\frac{k-1}{K}\right)}\sum_{f\in\mathcal{B}^{\ast}_{k}}\omega^{-1}_{f}L(1/2,F\otimes f)=\frac{L(1,F)K}{4}\widehat{W}(0)+O(K^{-1/4+\varepsilon}),$ (1.7)
where $\widehat{W}$ denotes the ordinary Fourier transform of $W$.
Consequently, as a by-product we can obtain a cute result for the non-vanishing problem of $L(1/2,F\otimes f)$ as follows:
@@
For any self-dual Maaß from $F$ on ${\mathrm{GL}}(3)$, there exists a newform $f\in\mathcal{B}^{\ast}_{k}$ with $k$ sufficiently large, such that $L(1/2,F\otimes f)\neq 0$.
It is instructive to see that the estimate in (1.6) should be the best-possible in essence, and is already able to beat what the classic spectral large sieve (see, for example, [12, Theorem 2]) together with the Cauchy-Schwarz inequality implies, which only provides a quannity of $O(K^{2+\varepsilon})$. This also reflects the significant effectiveness by further averaging over the parameter $k\sim K$, when it comes to investigating the subconvexity problem involving the weight aspect. We also refer the readers to the works of Luo [39] and Khan [27] for relevant heuristics. It, on the other hand, should be noticed that, thanks to the rational structure of the exponential factor in (4.7), we would succeed to achieve this strong Lindelöf average bound; see Remark 2.10 below for details. Besides, it is also believed that our method in proving Theorem 1.6 can be naturally adapted to produce simultaneous non-vanishing of ${\mathrm{GL}}(3)\times{\mathrm{GL}}(2)$ and ${\mathrm{GL}}(2)$ $L$-functions at the central point, and even subconvexity for $L(1/2,F\otimes f)$ or $L(1/2,f)L(1/2,F\otimes f)$ in the weight and the level (of $F$ or $f$) aspects. We are now working on this direction, and expect to report the progress in the future article.
The initial manoeuvre of our proofs of main results follows the path pioneered by Xiaoqing Li and Blomer, who looked into the average of first moment of $\rm GL(3)\times GL(2)$ $L$-function in a family. The strategies we pursued in the paper, however, can be understood as the “Kuznetsov-Voronoĭ-Reciprocity-Voronoĭ-Bilinear form” in proving Theorem 1.2 and “Kuznetsov-Exponential sums/Voronoĭ” in proving Theorems 1.4 & 1.7, respectively. In particular, for the former, this is reminiscent of an important work of Blomer and Khan, [6], in which they investigated the twisted moments of $\text{GL}(3)\times\text{GL}(2)$ and $\text{GL}(2)$ automorphic $L$-functions, and established the following reciprocity formula
@@
by the “Kuznetsov-Voronoĭ-Reciprocity-Voronoĭ-Kuznetsov” law. Among other things, one salient point is the application of the reciprocity law, which, as the Kloosterman sum of dimension one, keeps us going back to somewhere. However, when applying the Voronoĭ to the Kloosterman sum $S(n\ell,1;c)$ of dimension two, the inverse of the variable $\ell$, that is, the exponential factor like $S(n\overline{\ell},\sigma;c)$ for some less significant parameter $\sigma$, will be occurred; this will increase the dimensions of Kloosterman sums afterwards, if one were to resort to the Voronoĭ again. In this sense, we can barely expect the reciprocity formula for Hecke-Maaß forms of higher degrees $d\geq 4$ in the levels aspects [^2] Alternatively, one might explore the reciprocity by equipping with the Kuznetsov formulae in the level aspect for ${\rm{GL}}(n)$, $n\geq 3$; see, for example, [4] for descriptions on ${\rm{GL}}(3)$. However, observe that, it seems very tricky to establish the relation $S(m_{1}m_{2}\cdots m_{k}{c_{1}}\overline{\ell},1;c_{2})\leftrightsquigarrow S(n_{1}n_{2}\cdots n_{k}{c_{1}}\overline{q},1;c_{2})$ for the levels $\ell,q\geq 1$, where $m_{i},n_{i}$ are related to the summations of Fourier coefficients of ${\mathrm{GL}}(d_{i})$-forms, $d_{i}\geq 1$. It is still not sure that if a kind of reciprocity law involving the ${\mathrm{GL}}(3)$ level aspect can be achieved.. It seems that the possible symmetric relation like (1.8) for $\text{GL}(d)\times\text{GL}(2)$ involving the levels aspects is only for the $\text{GL}(3)\times\text{GL}(2)$ case (we refer the interested readers to Jana and Nunes’s recent work [25] for the generalization to $\rm{PGL}_{d+1}\times\rm{PGL}_{d-1}$). In contrast to the $\text{GL}(3)\times\text{GL}(2)$ case, we notice that, for the $\text{GL}(2)\times\text{GL}(1)$ case, Conrey [10] proved the conductors aspects reciprocity formula
@@
Let us now elucidate two of main obstacles in the present paper. It is remarkable that, by the current development of the Voronoĭ formulae for ${\mathrm{GL}}(3)$ in the level aspect, there is still a gap for arbitrary non-zero integer $c\in\mathbb{Z}$ (see Lemma 2.1 below). If the additive twist colludes with the level $N$ of the form $F$, namely, the parameters $c,N$ satisfy that $(N/(N,c),c)\neq 1$. For instance, one takes $N=q^{2}$ and $c=c^{\prime}q$ with $(c^{\prime},q)=1$ for any prime $q$. The ${{\mathrm{GL}}}(3)$-Voronoĭ formula in this case looks rather tricky. In principle, Corbett’s formula [9, Theorem 1.1] can cover this, but it requires a non-trivial analysis of the $p$-adic Bessel transforms, which, however, becomes more involved; see [9] for relevant details. We also refer the readers to the work of Booker, Milinovich and Ng [8], who worked out the Voronoĭ formula in the ${{\mathrm{GL}}}(2)$-case (if the additive twist colludes with the level). As an innovation in this paper, we are able to overcome this deadlock, which makes the hybrid subconvexity for ${\mathrm{GL}}(3)$ and ${\mathrm{GL}}(2)$ in the levels aspects come into being. The key point, however, is to invoke the reciprocity law as mentioned above. Indeed, with a suitable constraint for the parameter $\ell$ we are in the unramified case, when applying the Voronoĭ for the first time; then, one would be transferred to the ramified case after an implementation of the reciprocity law. This makes it possible for one to equip with the Voronoĭ again.
The second obstacle is the root number. It is well-known that the explicit determination of root numbers play a focal rôle when doing averages of central values of $L$-functions in different families. In this paper, we will give a detailed description of the root number $\varepsilon(F\otimes f)$ for any ${\mathrm{GL}}(3)$ Hecke-Maaß form $F$ with arbitrary level $N$ in the framework of the adelic interpretations, which is of independent interest; see §2.3. Part of descriptions in the special case where $F$ arises as a symmetric square lift from ${\mathrm{GL}}(2)$ has already been given by Holowinsky, Munshi and Qi [18]. As it can be seen later, the root number $\varepsilon(F\otimes f)$ depends only on the parameter $k$ and the form $F$. One of its factor $i^{k}$ will have a direct influence on the ‘harmonic’ averages of the Bessel function $J_{k-1}$ in considering the moments of $L$-values in the weight $k$ aspect; see Lemmas 2.8 and 2.9.
@@
Throughout the paper, $\varepsilon$ (resp. $\delta$) always denotes an arbitrarily small positive constant which may not be the same at each occurrence. $A\ll_{\varepsilon}B$ means that $|A|\ll X^{\varepsilon}|B|$, and $n\sim N$ means that $N<n\leq 2N$ for every $n\in\mathbb{Z}$. For any positive integers $m,n$, $(m,n)$ denotes the great common divisor of $m$ and $n$. $\mu$ is the Möbius function.